OpenStudy (anonymous):

Find cot theta if csc theta =-sqrt37/6 and tan theta>0

OpenStudy (freckles):

\[\sin^2(x)+\cos^2(x)=1 \\ \text{ divide both sides by } \sin^2(x) \\ 1+\cot^2(x)=\csc^2(x) \\ 1+(\cot(x))^2=(\csc(x))^2 \\ \]

OpenStudy (freckles):

replace csc(x) with -sqrt(37)/6 and then solve for cot(x)

OpenStudy (freckles):

and yes I changed your theta to x

OpenStudy (freckles):

keep in mind if tangent is positive then so is cotangent so you will take the positive answer

OpenStudy (anonymous):

Thank you!

OpenStudy (freckles):

np